how many 7" high x 16" long bricks will fit in a 72" wide x 120" high rectangle?
Answer by ikleyn(52908)

.
how many 7" high x 16" long bricks will fit in a 72" wide x 120" high rectangle?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

I can put 10 times  7" along the  72" side (10 bricks) 
      and  7 times 16" along the 120" side ( 7 bricks)

(the simplest and the most obvious placement).

In this way I will put 10*7 = 70 bricks.

The entire area of the original rectangle is 72*120 = 8640 sq. in.

The covered are will be                      70*112 = 7840 sq. in.

The difference (the uncovered area) will be            800 sq. in.

The area of a single brick is                  7*16 =  112 sq. in.

The uncovered area is equivalent to 7+ bricks.

And what?